### v1

@@ 1 Introduction
The celebrated VCG mechanism can be used to handle the incentives issue by inducing truthful reporting as dominant strategies. Unfortunately, in trade settings the VCG mechanism is not budget balanced but rather loses money, and thus is not feasible without an external source of subsidies. As the VCG mechanism is essentially the unique mechanism that maximizes gains from trade in dominant strategies, this implies the general impossibility of maximizing gains from trade in a budget balanced way in dominant strategies. In fact, [MS83] show that maximizing gains from trade without subsidies is impossible even in the much more relaxed sense of Bayesian incentive compatibility. [^6] And even if we only require interim individual rationality and weak budget balance. Furthermore, this holds even for the setting of a bilateral trade in which a single seller has one item to sell to a single buyer.
@@
Let us start with the simplest trade setting, that of bilateral trade. For that setting the learning problem is well understood. In a bilateral trade setting there is a single buyer and single seller with an item to sell, and their private values for the item are $v_{s}$ and $v_{b}$, respectively. These pairs of values $(v_{s},v_{b})$ are jointly distributed according to a distribution V. The goal is to find a simple mechanism with maximum GFT over all simple mechanisms. The starting point for addressing this type of question is a characterization of the family of possible simple mechanisms, and then optimize over that family. Luckily, simple mechanisms for bilateral trade have a very limited form: they are fixed price mechanisms. That is, any simple mechanism is defined by a fixed price $p$, and trade occurs, at price $p$, if $v_{s}<p<v_{b}$, and no trade happens if $v_{s}>p$ or $p>v_{b}$ (in case of ties, trade may or may not happen). It follows that the challenge in this case is to learn a price $p$ such that the fixed-price mechanism with price $p$ maximizes the gains from trade for V (over all such fixed-price mechanisms). Since this requires learning just a single parameter, $p$, this is indeed doable and, as implied by [Ces+21] [^7] This paper actually solves a harder problem, in a regret-minimization setting, as well as analyzes several related models. See Section 1.1 for details., choosing $O(\varepsilon^{-1})$ sample points and picking the sampled value [^8] It is interesting to note, though, that it does not suffice to choose among some predetermined $\delta$-grid of values. that yields highest GFT on the sample gives, with high probability, a mechanism that is optimal up to an additive $\varepsilon$ on the real, yet unknown, distribution V. Significantly, this holds for any joint distribution V, and does not require the buyer’s and seller’s values to be independent.
@@
Given the characterization above, it is natural to attempt learning a (nearly) optimal simple mechanism by optimizing over all mechanisms $M_{f_{1},f_{2}}$, over all pairs of compatible functions $(f_{1},f_{2})$. Unfortunately, this class turns out to have infinite “dimension”, and in fact, we show that no low-dimension subclass suffices for even approximating the gains from trade. The impossibility result is very general and shows that it is even impossible to learn a “general” mechanism in the sense of [MS83] that need only be Bayesian incentive compatible, interim individually rational, and weakly budget balanced.
@@
The simple mechanism that extracts the first-best GFT can be viewed as a perfect over-fitting of the mechanism to the $k$ triplets in $S$. This simple lemma implies the theorem, since for any finite number of samples $t$ it is impossible to distinguish between the following two cases: 1) the $t$ samples are coming from the uniform distribution over $[0,1]^{3}$, for which there is a gap between the first best and what a strategic mechanism can achieve, and 2) the $t$ samples are from $U(S)$ for a “random” generic set $S$ of size $k>>t$, for which there is no such gap. This impossibility is due to the fact that for large enough $k$, it is unlikely that we sample any element of $S$ more than once. [^11] Note that for the bilateral-trade setting an appropriate adjustment of the claim of Lemma 1.3 is not true (consistently with the fact that Theorem 1 does not hold for bilateral trade). That is, consider the set $S$ with two samples $(s,b)$ taking the realizations $(0,2)$ and $(5,7)$ (note that this set of values is generic). As any simple mechanism for bilateral trade is a fixed-price mechanism, the maximum expected GFT of any simple mechanism on a uniform distribution over these two samples is $1$, while the first-best GFT is $2$.
@@
Figure 4: Optimal pair for $v_{s}\sim U[0,1]$, $v_{1}\sim\begin{cases}U[0,\frac{1}{2}],&\frac{1}{4}\\ U[\frac{1}{2},\frac{3}{4}],&\frac{3}{4}\end{cases}$, $v_{2}\sim\begin{cases}U[0,\frac{1}{4}],&\frac{3}{4}\\ U[\frac{1}{2},1],&\frac{1}{4}\end{cases}$. Note that the functions are rather complicated.
@@ 1.1 Additional related work
Prior work [Ces+21] has studied the bilateral trade problem in a regret-minimization framework over rounds of seller/buyer interactions, with no prior knowledge on the private seller/buyer valuations. They showed that learning a close to optimal simple mechanism can be done in some challenging online scenarios (but provably not in some even more challenging ones). We study the easier problem of (offline) learning from samples, but the setting of 1-seller 2-buyers and beyond (while they study bilateral trade).
The problem of gains-from-trade maximization in trade settings has been extensively studied. Myerson and Satterthwaite [MS83] show that for bilateral trade settings, the second-best mechanism has strictly lower GFT than the first best. The second-best mechanism is the mechanism with the highest GFT subject to being Bayesian Incentive-Compatible (BIC), Interim Individually Rational (IIR), and ex-ante Weakly Budget Balanced (WBB). In the double auctions setting, McAfee[McA92] has shown that there is a DSIC mechanisms that is IIR and WBB with GFT that is at least $(1-1/q)$ fraction of the first best, when $q$ is the size of the efficient trade. This result was extended to other trade settings, e.g., to spatially-distributed markets [BNP09]. Note that the approximation provided by this mechanism to the first best is 00 when $q=1$, and thus this mechanism gives no approximation in small markets, and in particular, in single-seller settings (e.g., bilateral trade and the 1-seller 2-buyer setting we focus on). Recently, it was shown that for bilateral trade (and for double auctions), the multiplicative gap between the first and second best is only a constant factor [Den+22].
For double auctions, [Bab+18] has presented mechanisms that are constant approximation to the second best, and are also asymptotically efficient. In [BGG20], a Bulow-Klemperer-style result for GFT maximization was presented to the double auctions setting, showing that with IID agents, a variant of the McAfee[McA92] mechanism with an additional buyer ensures at least as much expected GFT as the first best of the original setting.
A long line of research has studied constant-fraction approximation to the welfare in bilateral trade [BD21, CW23, KPV22] and more general trade settings [Col+20].
@@ 2.1 Simple Mechanisms
As usual, we focus on mechanisms that are normalized (losers pay zero):
@@
A celebrated result by Myerson [Mye81] implies a characterization of normalized DSIC mechanisms in a single-parameter domain. We will use a simple statement of this result, based on Theorem 9.36 in [Nis+07], as it applies for our domain:
@@ 2.2 Gains From Trade
In this paper we focus on optimizing the gains form trade (GFT). Given a joint prior distribution V over triplet of values $\textit{{v}}=(v_{s},v_{1},v_{2})\in\mathbb{R}_{\geq 0}^{3}$, the unconstrained optimal GFT (first best), denoted by $\textup{OPT}(\boldsymbol{V})$, is defined to be
@@ 3.1 Associated Functions
[…]
We argue that these functions are indeed well defined (e.g., $f_{1}(v_{2})$ does the depend on the choice of $v^{\prime}_{s}$ and $v^{\prime}_{1}$). As $M$ is a DSIC mechanism, by the Taxation Principle, the payment of each buyer when winning cannot be affected by his own report. Additionally, as the mechanism is SBB and normalized, when there is a trade, the seller receives the payment of the trading buyer, and thus the trading buyer’s payment cannot depend on the report of the seller. Therefore, there is a single parameter function that defines the payment in such a trade, and it can only depend on the non-trading buyer’s value.
@@
Intuitively, the functions $f_{1}$ and $f_{2}$ represent the critical values for the buyers to trade, so they can never imply that both buyers trade concurrently. The compatibility condition does exactly that - it disallows functions that imply that both buyers are each bidding strictly above his critical value.
@@ 3.2 The Best Mechanism for Compatible Functions
The mechanism described in Definition 3.4 preforms a trade whenever a trade increases the GFT, and it break ties in favor of higher GFT, and thus it maximizes GFT among all mechanisms that are associated with the pair of functions $(f_{1},f_{2})$:
@@ 4 Impossibility of Learning for General Distributions
In this section we show that for unrestricted joint distributions (when values might be correlated), it is impossible to learn an approximately optimal simple mechanism, even when the support is on $[0,1]^{3}$. To begin, we show that for every finite joint distribution that satisfies a mild condition (“generic support)”, there is a simple mechanism that achieves the maximum possible GFT (that is, there is a simple mechanism that obtains the first-best GFT). Subsequently, we leverage this finding to demonstrate scenarios where unrestricted joint distributions pose a challenge to the feasibility of learning approximately optimal mechanisms.
@@ 4.1 First-best GFT by a Simple Mechanism
Formally, to be able to extract all GFT it is enough that the distribution is supported on a generic set of vectors. A set $C=\{(v_{1}^{1},v_{2}^{1},...,v_{d}^{1}),...,(v_{1}^{n},v_{2}^{n},...,v_{d}^{n})\}\subseteq\mathbb{R}^{d}_{\geq 0}$ of $n$ vectors (of length $d$) is called generic, if there are no two values that are identical (for all $i_{1},i_{2}\in[d]$ and $j_{1},j_{2}\in[n]$ it holds that $v_{i_{1}}^{j_{1}}\neq v_{i_{2}}^{j_{2}}$ unless $i_{1}=i_{2}$ and $j_{1}=j_{2}$). [^15] For our proof to hold it is sufficient that a weaker condition holds: no value repeats more than once in any fixed coordinate $i\in[d]$.
@@
Let $C=\{(v_{s}^{1},v_{1}^{1},v_{2}^{1}),...,(v_{s}^{n},v_{1}^{n},v_{2}^{n})\}$ be a generic set of vectors representing the support of a finite joint distribution V for values in the 1-seller 2-buyer setting. We define the functions $f_{1}^{C},f_{2}^{C}$ as follows:
[…]
@@
These functions capitalize on the knowledge encoded within $C$ to establish distinct prices for each triplet of values. They achieve this by consistently setting the price as the higher buyer’s value while setting the other’s to $\infty$. The price for each buyer is set using a condition only on the report of the other buyer, as necessary for a DSIC mechanism. This approach renders the higher buyer indifferent and thus prompts acceptance of the trade, whereas the lower buyer is either not offered the trade or is presented with an offer at $\infty$ (equivalent to non-offer). Additionally, the functions are compatible because for every $k\in[n]$, if the function $f_{b}^{C}(v_{-b}^{k})$ is set to be some value that is not $\infty$, then either $f_{-b}^{C}(v_{b}^{k})=\infty$ or $f_{b}^{C}(v_{-b}^{k})=v_{b}^{k}=v_{-b}^{k}=f_{-b}^{C}(v_{b}^{k})$ and both cases do no violate compatibility. Hence, the mechanism $M_{f_{1}^{C},f_{2}^{C}}$ is straightforwardly a well-defined simple mechanism, a fact that can also be directly deduced by Lemma 3.5.
@@
Let $C$ be a finite generic set, and let V be a probability distribution over that set. Then the GFT of the simple mechanism $M_{f_{1}^{C},f_{2}^{C}}$ is the unconstrained optimum (first best): $\textup{GFT}(M_{f_{1}^{C},f_{2}^{C}},\textbf{V})=\textup{OPT}(\boldsymbol{V})$.
@@ 4.2 Impossibility of Learning for General Distributions
We will be applying this lemma to cases where not only is $\textup{OPT}_{\textup{simple}}(\boldsymbol{V})$ smaller than $\textup{OPT}(\boldsymbol{V})$, but even the GFT of any Bayesian incentive-compatible and weakly budget-balanced mechanism, as studied by [MS83], is smaller. So for explicitness let us denote the GFT of the best mechanism from the general class studied by [MS83] (the “second best” there) by $\textup{OPT}_{\textup{BIC}}(\boldsymbol{V})$. Note that $\textup{OPT}_{\textup{simple}}(\boldsymbol{V})\leq\textup{OPT}_{\textup{BIC}}(\boldsymbol{V})$ and whenever $\textup{OPT}(\boldsymbol{V})=\textup{OPT}_{\textup{simple}}(\boldsymbol{V})$ then also $\textup{OPT}(\boldsymbol{V})=\textup{OPT}_{\textup{BIC}}(\boldsymbol{V})$. We now start with some uncorrelated distribution where the first-best $\textup{OPT}(\boldsymbol{V})$ is strictly larger than the second-best $\textup{OPT}_{\textup{BIC}}(\boldsymbol{V})$ (and thus certainly also larger than $\textup{OPT}_{\textup{simple}}(\boldsymbol{V})$) [MS83]. Specifically, $v_{s},v_{1}$ and $v_{2}$ are independently sampled, each distributed uniformly on $[0,1]$. [^16] Actually, [MS83] does not explicitly discuss cases apart from bilateral trade. However, for 1-seller 2-buyer uniform distribution over $[0,1]^{3}$, consider the subcase when $v_{s},v_{1}\geq\frac{1}{2}$ and $v_{2}<\frac{1}{2}$ (which happens with a constant probability). Since no trade with buyer 2 is possible, and since buyer $2$ value gives no useful information due to independence, we revert to a bilateral trade setting where [MS83] showed a gap - and so there must also be a gap for uniform distribution over $[0,1]^{3}$.
@@
On the other hand, since we specifically started with a distribution for which there is a gap between the first best and second best (the uniform distribution over $[0,1]^{3}$), the GFT of any simple mechanism is smaller than the first-best by some constant [MS83]. Assuming that we have an algorithm that learns, using $t$ samples, a simple mechanism that (almost) maximizes gains from trade, then the previous lemma implies that we can distinguish between the (correlated) distribution $U_{S}$, for a random $S$, and the uniform distribution over $([0,1]^{3})^{t}$. But that, as we will show, is impossible, as these distributions are statistically close to each other.
@@
Combining these two lemmata we get the result showing that optimal simple mechanisms cannot be learned, and in fact it is even impossible to learn a general mechanism (Bayesian incentive-compatible and weakly budget-balanced in the sense of [MS83]) that approximates the GFT of the best simple mechanism.
@@
Assume that such a learning algorithm exists for $c>0$ small enough so that $6c$ is smaller than the gap between the first-best and the second best GFT of the uniform distribution over $[0,1]^{3}$, as ensured by [Mye81] (see Footnote 16). Lemma 4.3 shows that the algorithm may be used to distinguish between a uniform distribution on any generic set for which, by Lemma 4.2, $\textup{OPT}_{\textup{simple}}(\boldsymbol{V})=\textup{OPT}(\boldsymbol{V})$, and the uniform distribution on $[0,1]^{3}$ for which $\textup{OPT}_{\textup{simple}}(\boldsymbol{V})\leq\textup{OPT}_{\textup{BIC}}(\boldsymbol{V})\leq\textup{OPT}(\boldsymbol{V})-6c$. Since the distribution $U_{S}$ from the previous lemma is just the average over all possible sets $S$ of size $T$ (which are generic with probability 1) of the uniform distribution of $t$ samples from the set $S$, the algorithm (which accepts $t$ samples from $[0,1]^{3}$) also separates between $U_{S}$ and the uniform distribution on $([0,1]^{3})^{t}$. But this yields a contradiction as Lemma 4.4 states that no algorithm can achieve such a separation with high probability.

### v2

@@ 1 Introduction
The celebrated VCG mechanism can be used to handle the incentives issue by inducing truthful reporting as dominant strategies. Unfortunately, in trade settings the VCG mechanism is not budget balanced but rather loses money, and thus is not feasible without an external source of subsidies. As the VCG mechanism is essentially the unique mechanism that maximizes gains from trade in dominant strategies, this implies the general impossibility of maximizing gains from trade in a budget balanced way in dominant strategies. In fact, [myerson1983efficient] show that maximizing gains from trade without subsidies is impossible even in the much more relaxed sense of Bayesian incentive compatibility. [^6] And even if we only require interim individual rationality and weak budget balance. Furthermore, this holds even for the setting of a bilateral trade in which a single seller has one item to sell to a single buyer.
@@
Let us start with the simplest trade setting, that of bilateral trade. For that setting the learning problem is well understood. In a bilateral trade setting there is a single buyer and single seller with an item to sell, and their private values for the item are $v_{s}$ and $v_{b}$, respectively. These pairs of values $(v_{s},v_{b})$ are jointly distributed according to a distribution V. The goal is to find a simple mechanism with maximum GFT over all simple mechanisms. The starting point for addressing this type of question is a characterization of the family of possible simple mechanisms, and then optimize over that family. Luckily, simple mechanisms for bilateral trade have a very limited form: they are fixed price mechanisms. That is, any simple mechanism is defined by a fixed price $p$, and trade occurs, at price $p$, if $v_{s}<p<v_{b}$, and no trade happens if $v_{s}>p$ or $p>v_{b}$ (in case of ties, trade may or may not happen). It follows that the challenge in this case is to learn a price $p$ such that the fixed-price mechanism with price $p$ maximizes the gains from trade for V (over all such fixed-price mechanisms). Since this requires learning just a single parameter, $p$, this is indeed doable and, as implied by [Cesa-BianchiCCF21] [^7] This paper actually solves a harder problem, in a regret-minimization setting, as well as analyzes several related models. See Section 1.1 for details., choosing $O(\varepsilon^{-1})$ sample points and picking the sampled value [^8] It is interesting to note, though, that it does not suffice to choose among some predetermined $\delta$-grid of values. that yields highest GFT on the sample gives, with high probability, a mechanism that is optimal up to an additive $\varepsilon$ on the real, yet unknown, distribution V. Significantly, this holds for any joint distribution V, and does not require the buyer’s and seller’s values to be independent.
@@
Given the characterization above, it is natural to attempt learning a (nearly) optimal simple mechanism by optimizing over all mechanisms $M_{f_{1},f_{2}}$, over all pairs of compatible functions $(f_{1},f_{2})$. Unfortunately, this class turns out to have infinite “dimension”, and in fact, we show that no low-dimension subclass suffices for even approximating the gains from trade. The impossibility result is very general and shows that it is even impossible to learn a “general” mechanism in the sense of [myerson1983efficient] that need only be Bayesian incentive compatible, interim individually rational, and weakly budget balanced.
@@
The simple mechanism that extracts the first-best GFT can be viewed as a perfect over-fitting of the mechanism to the $k$ triplets in $S$. number of samples $t$ it is impossible to distinguish between the following two cases: We show that this simple lemma implies the theorem, by proving that for any finite number of samples $t$, it is impossible to distinguish between the following two cases: 1) the $t$ samples are coming from the uniform distribution over $[0,1]^{3}$, for which there is a gap between the first best and what a strategic mechanism can achieve, and 2) the $t$ samples are from $U(S)$ for a “random” generic set $S$ of size $k>>t$, for which there is no such gap. This impossibility is due to the fact that for large enough $k$, it is unlikely that we sample any element of $S$ more than once. [^11] Note that for the bilateral-trade setting an appropriate adjustment of the claim of Lemma 1.3 is not true (consistently with the fact that Theorem 1 does not hold for bilateral trade). That is, consider the set $S$ with two samples $(s,b)$ taking the realizations $(0,2)$ and $(5,7)$ (note that this set of values is generic). As any simple mechanism for bilateral trade is a fixed-price mechanism, the maximum expected GFT of any simple mechanism on a uniform distribution over these two samples is $1$, while the first-best GFT is $2$.
@@
Figure 4: Optimal pair for $v_{s}\sim U[0,1]$,
$v_{1}\sim\begin{cases}U[0,\frac{1}{2}],&\frac{1}{4}\\ U[\frac{1}{2},\frac{3}{4}],&\frac{3}{4}\end{cases}$, $v_{2}\sim\begin{cases}U[0,\frac{1}{4}],&\frac{3}{4}\\ U[\frac{1}{2},1],&\frac{1}{4}\end{cases}$.
Note that the functions are rather complicated.
@@ 1.1 Additional related work
Prior work [Cesa-BianchiCCF21] has studied the bilateral trade problem in a regret-minimization framework over rounds of seller/buyer interactions, with no prior knowledge on the private seller/buyer valuations. They showed that learning a close to optimal simple mechanism can be done in some challenging online scenarios (but provably not in some even more challenging ones). We study the easier problem of (offline) learning from samples, but the setting of 1-seller 2-buyers and beyond (while they study bilateral trade).
The problem of gains-from-trade maximization in trade settings has been extensively studied. Myerson and Satterthwaite [myerson1983efficient] show that for bilateral trade settings, the second-best mechanism has strictly lower GFT than the first best. The second-best mechanism is the mechanism with the highest GFT subject to being Bayesian Incentive-Compatible (BIC), Interim Individually Rational (IIR), and ex-ante Weakly Budget Balanced (WBB). In the double auctions setting, McAfee[McAfee92] has shown that there is a DSIC mechanisms that is IIR and WBB with GFT that is at least $(1-1/q)$ fraction of the first best, when $q$ is the size of the efficient trade. This result was extended to other trade settings, e.g., to spatially-distributed markets [BNP09]. Note that the approximation provided by this mechanism to the first best is 00 when $q=1$, and thus this mechanism gives no approximation in small markets, and in particular, in single-seller settings (e.g., bilateral trade and the 1-seller 2-buyer setting we focus on). Recently, it was shown that for bilateral trade (and for double auctions), the multiplicative gap between the first and second best is only a constant factor [DengMSW22].
For double auctions, [BCGZ18] has presented mechanisms that are constant approximation to the second best, and are also asymptotically efficient. In [BGG2020], a Bulow-Klemperer-style result for GFT maximization was presented to the double auctions setting, showing that with IID agents, a variant of the McAfee[McAfee92] mechanism with an additional buyer ensures at least as much expected GFT as the first best of the original setting.
A long line of research has studied constant-fraction approximation to the welfare in bilateral trade [BlumrosenD21, CW23, KPV2022] and more general trade settings [CGKLTS2020].
@@ 2.1 Simple Mechanisms
As usual, we focus on mechanisms that are normalized (losers pay zero) [^15] Note that this property does not necessarily hold for all ex-post IR and SBB mechanisms. E.g., a mechanism that offers a trade $b\rightarrow 1$ with $p_{1}=1,p_{2}=-1/2,p_{s}=1/2$. If the agents agree to the trade, Buyer 2 is a loser but his price is not 0. :
@@
A celebrated result by Myerson [myerson1981optimal] implies a characterization of normalized DSIC mechanisms in a single-parameter domain. We will use a simple statement of this result, based on Theorem 9.36 in [agt], as it applies for our domain:
@@ 2.2 Gains From Trade
In this paper we focus on optimizing the gains form trade (GFT). Given a joint prior distribution V over triplet of values $\textit{{v}}=(v_{s},v_{1},v_{2})\in\mathbb{R}_{\geq 0}^{3}$, the unconstrained optimal GFT (first best), denoted by $\textup{OPT}(\boldsymbol{V})$, is defined to be:
@@ 3.1 Associated Functions
$f_{1}(v_{2})=\begin{cases}p_{1}(v^{\prime}_{s},v^{\prime}_{1},v_{2}),&\exists v^{\prime}_{s},v^{\prime}_{1}:A(v^{\prime}_{s},v^{\prime}_{1},v_{2})=b_{1}\\ \infty,&\text{otherwise}\end{cases}$
$f_{2}(v_{1})=\begin{cases}p_{2}(v^{\prime}_{s},v_{1},v^{\prime}_{2}),&\exists v^{\prime}_{s},v^{\prime}_{2}:A(v^{\prime}_{s},v_{1},v^{\prime}_{2})=b_{2}\\ \infty,&\text{otherwise}\end{cases}$
We argue that these functions are indeed well defined (e.g., $f_{1}(v_{2})$ does not depend on the choice of $v^{\prime}_{s}$ and $v^{\prime}_{1}$). As $M$ is a DSIC mechanism, by the Taxation Principle, the payment of each buyer when winning cannot be affected by his own report. Additionally, as the mechanism is SBB and normalized, when there is a trade, the seller receives the payment of the trading buyer, and thus the trading buyer’s payment cannot depend on the report of the seller. Therefore, there is a single parameter function that defines the payment in such a trade, and it can only depend on the non-trading buyer’s value.
@@
Intuitively, the functions $f_{1}$ and $f_{2}$ represent the critical values for the buyers to trade, so they can never imply that both buyers trade concurrently. The compatibility (“non-crossing”) condition does exactly that - it disallows functions that imply that both buyers are each bidding strictly above his critical value.
@@ 3.2 The Best Mechanism for Compatible Functions
The mechanism described in Definition 3.4 performs a trade whenever a trade increases the GFT, and it break ties in favor of higher GFT, and thus it maximizes GFT among all mechanisms that are associated with the pair of functions $(f_{1},f_{2})$:
@@ 4 Impossibility of Learning for General Distributions
In this section we show that for unrestricted joint distributions (when values might be correlated), it is impossible to learn an approximately optimal simple mechanism, even when the support is on $[0,1]^{3}$. To begin, we show that for every finite joint distribution that satisfies a mild condition (“generic support”), there is a simple mechanism that achieves the maximum possible GFT (that is, there is a simple mechanism that obtains the first-best GFT). Subsequently, we leverage this finding to demonstrate scenarios where unrestricted joint distributions pose a challenge to the feasibility of learning approximately optimal mechanisms.
@@ 4.1 First-best GFT by a Simple Mechanism
Formally, to be able to extract all GFT it is enough that the distribution is supported on a generic set of vectors. A set $C=\{(v_{1}^{1},v_{2}^{1},\ldots,v_{d}^{1}),\ldots,(v_{1}^{n},v_{2}^{n},\ldots,v_{d}^{n})\}\subseteq\mathbb{R}^{d}_{\geq 0}$ of $n$ vectors (of length $d$) is called generic, if there are no two values that are identical (for all $i_{1},i_{2}\in[d]$ and $j_{1},j_{2}\in[n]$ it holds that $v_{i_{1}}^{j_{1}}\neq v_{i_{2}}^{j_{2}}$ unless $i_{1}=i_{2}$ and $j_{1}=j_{2}$). [^16] For our proof to hold it is sufficient that a weaker condition holds: no value repeats more than once in any fixed coordinate $i\in[d]$.
@@
Let $C=\{(v_{s}^{1},v_{1}^{1},v_{2}^{1}),\ldots,(v_{s}^{n},v_{1}^{n},v_{2}^{n})\}$ be a generic set of vectors representing the support of a finite joint distribution V for values in the 1-seller 2-buyer setting. We define the functions $f_{1}^{C},f_{2}^{C}$ as follows:
$\forall b\in\{b_{1},b_{2}\}:f_{b}^{C}(x)=\begin{cases}v_{b}^{k}&\exists k:(v_{s}^{k},v_{1}^{k},v_{2}^{k})\in C\ s.t.\ x=v_{-b}^{k}\leq v_{b}^{k}\\ \infty&\text{otherwise}\end{cases}$
@@
These functions capitalize on the knowledge encoded within $C$ to establish distinct prices for each triplet of values. They achieve this by consistently setting the price as the higher buyer’s value while setting the other’s to $\infty$. The price for each buyer is set using a condition dependent only on the report of the other buyer, as necessary for a DSIC mechanism. This approach renders the higher buyer indifferent and thus prompts acceptance of the trade, whereas the lower buyer is either not offered the trade or is presented with an offer at $\infty$ (equivalent to non-offer). Additionally, the functions are compatible because for every $k\in[n]$, if the function $f_{b}^{C}(v_{-b}^{k})$ is set to be some value that is not $\infty$, then either $f_{-b}^{C}(v_{b}^{k})=\infty$ or $f_{b}^{C}(v_{-b}^{k})=v_{b}^{k}=v_{-b}^{k}=f_{-b}^{C}(v_{b}^{k})$ and both cases do no violate compatibility. Hence, the mechanism $M_{f_{1}^{C},f_{2}^{C}}$ is straightforwardly a well-defined simple mechanism, a fact that can also be directly deduced by Lemma 3.5.
@@
Let $C$ be a finite generic set, and let V be a probability distribution over $C$. Then the GFT of the simple mechanism $M_{f_{1}^{C},f_{2}^{C}}$ is the unconstrained optimum (first best): $\textup{GFT}(M_{f_{1}^{C},f_{2}^{C}},\textbf{V})=\textup{OPT}(\boldsymbol{V})$.
@@ 4.2 Impossibility of Learning for General Distributions
We will be applying this lemma to cases where not only is $\textup{OPT}_{\textup{simple}}(\boldsymbol{V})$ smaller than $\textup{OPT}(\boldsymbol{V})$, but even the GFT of any Bayesian incentive-compatible and weakly budget-balanced mechanism, as studied by [myerson1983efficient], is smaller. So for explicitness let us denote the GFT of the best mechanism from the general class studied by [myerson1983efficient] (the “second best” there) by $\textup{OPT}_{\textup{BIC}}(\boldsymbol{V})$. Note that $\textup{OPT}_{\textup{simple}}(\boldsymbol{V})\leq\textup{OPT}_{\textup{BIC}}(\boldsymbol{V})$ and whenever $\textup{OPT}(\boldsymbol{V})=\textup{OPT}_{\textup{simple}}(\boldsymbol{V})$ then also $\textup{OPT}(\boldsymbol{V})=\textup{OPT}_{\textup{BIC}}(\boldsymbol{V})$. We now start with some uncorrelated distribution where the first-best $\textup{OPT}(\boldsymbol{V})$ is strictly larger than the second-best $\textup{OPT}_{\textup{BIC}}(\boldsymbol{V})$ (and thus certainly also larger than $\textup{OPT}_{\textup{simple}}(\boldsymbol{V})$) [myerson1983efficient]. Specifically, $v_{s},v_{1}$ and $v_{2}$ are independently sampled, each distributed uniformly on $[0,1]$. [^17] Actually, [myerson1983efficient] does not explicitly discuss cases apart from bilateral trade. However, for 1-seller 2-buyer uniform distribution over $[0,1]^{3}$, consider the subcase when $v_{s},v_{1}\geq\frac{1}{2}$ and $v_{2}<\frac{1}{2}$ (which happens with a constant probability). Since no trade with buyer 2 is possible, and since buyer $2$ value gives no useful information due to independence, we revert to a bilateral trade setting where [myerson1983efficient] showed a gap - and so there must also be a gap for uniform distribution over $[0,1]^{3}$.
@@
On the other hand, since we specifically started with a distribution for which there is a gap between the first best and second best (the uniform distribution over $[0,1]^{3}$), the GFT of any simple mechanism is smaller than the first-best by some constant [myerson1983efficient]. Assuming that we have an algorithm that learns, using $t$ samples, a simple mechanism that (almost) maximizes gains from trade, then the previous lemma implies that we can distinguish between the (correlated) distribution $U_{S}$, for a random $S$, and the uniform distribution over $([0,1]^{3})^{t}$. But that, as we will show, is impossible, as these distributions are statistically close to each other.
@@
Combining these two lemmata we get the result showing that optimal simple mechanisms cannot be learned, and in fact it is even impossible to learn a general mechanism (Bayesian incentive-compatible and weakly budget-balanced in the sense of [myerson1983efficient]) that approximates the GFT of the best simple mechanism.
@@
Assume that such a learning algorithm exists for $c>0$ small enough so that $6c$ is smaller than the gap between the first-best and the second best GFT of the uniform distribution over $[0,1]^{3}$, as ensured by [myerson1981optimal] (see Footnote 17). Lemma 4.3 shows that the algorithm may be used to distinguish between a uniform distribution on any generic set for which, by Lemma 4.2, $\textup{OPT}_{\textup{simple}}(\boldsymbol{V})=\textup{OPT}(\boldsymbol{V})$, and the uniform distribution on $[0,1]^{3}$ for which $\textup{OPT}_{\textup{simple}}(\boldsymbol{V})\leq\textup{OPT}_{\textup{BIC}}(\boldsymbol{V})\leq\textup{OPT}(\boldsymbol{V})-6c$. Since the distribution $U_{S}$ from the previous lemma is just the average over all possible sets $S$ of size $T$ (which are generic with probability 1) of the uniform distribution of $t$ samples from the set $S$, the algorithm (which accepts $t$ samples from $[0,1]^{3}$) also separates between $U_{S}$ and the uniform distribution on $([0,1]^{3})^{t}$. But this yields a contradiction as Lemma 4.4 states that no algorithm can achieve such a separation with high probability.